13. Rainbow owns the Brite Color Crayon Co. Blocks of colore d wax, which can be made into 100 crayons, cost $7.00 each. Rainbow estin testhe overhead expenses are 25% of the selling price, which is $2.50 for a pack of twelve crayons. What is the net profit rate per crayon. Hint: Start by calculating the supply price and selling price for a single crayon.
Answer #1

A block of colored wax, which makes 100 crayons cost $ 7.00.

=> Purchase price of 100 crayons = $ 7.00

=> Purchase price per crayon = $ 7.00 / 100 = $ 0.07

Selling Price for a pack of 12 Crayons = $ 2.50

Selling Price per Crayon = $ 2.50/ 12 = $ 0.2083

Overhead Expenses per crayon = 25% of Selling Price for a Single Crayon = 25% of $ 0.2083

= $ 0.0521

Net Profit per crayon = Selling Price per crayon - Purchase price per crayon - Overhead expenses per crayon

= $ (0.2083 - 0.07 - 0.0521) = $ 0.08625

Net Profit Rate = (Net Profit per crayon / Selling Price per crayon)*100

= (0.08625 / 0.2083)*100 = 41.41%

So, Net profit Rate per crayon = 41.41%